OpenStudy (anonymous):

The local library asked Reese's art class to paint a mural on a wall of the library. For the background, they painted a large square and used enough paint to cover 85ft squared . How long is one side of the square rounded to the nearest whole number?

OpenStudy (anonymous):

The choices are 6 feet 7 feet 8 feet 9 feet

OpenStudy (anonymous):

The area is 85 squared feet. It is a square. Area is equal to Length * Width. Because it's a square, you can simplify that to Area = Length^2 In other words, just square root both sides to find the length.

OpenStudy (anonymous):

i dont have any of the sides.. all i have is 85^2

OpenStudy (anonymous):

Sorry, I meant square root the area. Because: Area = Length^2 And you have the area.

Directrix (directrix):

Given that the square was painted with 85 sq feet of paint, think of 85 as the area of the square. Because the area of a square is (s^2), then 85 = s^2 and s = sqrt(85) = 9.2195. To the nearest whole number the side of the square is 9 feet. --> answer